# Robust Online Conformal Prediction under Uniform Label Noise

### v2

@@ Appendix C Proof for Section 3
Appendix D Proof for Section 4
D.1 Proof for Proposition 4.1
Lemma D.1 (Lemma 1 in Angelopoulos et al. (2024)).
$-\alpha\cdot\max_{1\leq t\leq T-1}\eta_{t}\leq\hat{\tau}_{T}\leq 1+(1-\alpha)\cdot\max_{1\leq t\leq T-1}\eta_{t}.$ (10)
Proposition D.2 (Restatement of Proposition 4.1).
$\alpha-\frac{1}{T}\sum_{t=1}^{T}\mathds{1}\left\{Y_{t}\notin\mathcal{C}_{t}(X_{t})\right\}\leq$ $\frac{2-\epsilon}{1-\epsilon}\sqrt{2\log\left(\frac{4}{\delta}\right)}\cdot\frac{1}{\sqrt{T}}+\frac{1+\underset{1\leq t\leq T-1}{\max}\eta_{t}}{1-\epsilon}\sum_{t=1}^{T}\left|\eta_{t}^{-1}-\eta_{t-1}^{-1}\right|\cdot\frac{1}{T}+$
$\frac{\epsilon}{1-\epsilon}\cdot\frac{1}{T}\sum_{t=1}^{T}\left((1-\alpha)-\frac{1}{K}\mathbb{E}\left[\mathcal{C}_{t}(X_{t})\right]\right).$
Proof.
Consider the gradient of clean pinball loss:
Part (a): Lemma G.3 gives us that
$\mathbb{E}_{S}\left[\nabla_{\hat{\tau}_{t}}l_{1-\alpha}(\hat{\tau}_{t},S_{t})\right]$ =$=\mathbb{P}\left\{S_{t}\leq\hat{\tau}_{t}\right\}-(1-\alpha)$
=$\overset{(a)}{=}\left(\frac{1}{1-\epsilon}\mathbb{P}\left\{\tilde{S}_{t}\leq\hat{\tau}_{t}\right\}-\frac{\epsilon}{K(1-\epsilon)}\sum_{y=1}^{K}\mathbb{P}\left\{S_{t,y}\leq\hat{\tau}_{t}\right\}\right)-(1-\alpha)$
where (a) comes from Lemma G.1. In addition, Lemma G.3 implies that
Thus, we can derive an upper bound for (b) as follows
$\leq\frac{1}{1-\epsilon}\left|\sum_{t=1}^{T}\mathbb{E}_{\tilde{S}}\left[\nabla_{\hat{\tau}_{t}}l_{1-\alpha}(\hat{\tau}_{t},\tilde{S})\right]-\sum_{t=1}^{T}\nabla_{\hat{\tau}_{t}}l_{1-\alpha}(\hat{\tau}_{t},\tilde{S})\right|+\frac{1}{1-\epsilon}\left|\sum_{t=1}^{T}\nabla_{\hat{\tau}_{t}}l_{1-\alpha}(\hat{\tau}_{t},\tilde{S})\right|+\frac{\epsilon}{1-\epsilon}\sum_{t=1}^{T}\left((1-\alpha)-\frac{1}{K}\mathbb{E}\left[\mathcal{C}_{t}(X_{t})\right]\right)$
Then, we derive an upper bound for (c). Define $\eta_{0}^{-1}=0$:
$\left|\sum_{t=1}^{T}\nabla_{\hat{\tau}_{t}}l_{1-\alpha}(\hat{\tau}_{t},S_{t})\right|$ $=\left|\sum_{t=1}^{T}\eta_{t}^{-1}\cdot\left(\eta_{t}\cdot\nabla_{\hat{\tau}_{t}}l_{1-\alpha}(\hat{\tau}_{t},S_{t})\right)\right|$
$=\left|\sum_{t=1}^{T}(\eta_{t}^{-1}-\eta_{t-1}^{-1})\cdot\left(\sum_{s=t}^{T}\eta_{s}\cdot\nabla_{\hat{\tau}_{s}}l_{1-\alpha}(\hat{\tau}_{s},S_{s})\right)\right|$
=$\overset{(a)}{=}\left|\sum_{t=1}^{T}(\eta_{t}^{-1}-\eta_{t-1}^{-1})\cdot(\hat{\tau}_{T}-\hat{\tau}_{t})\right|$
$\overset{(c)}{\leq}(1+\max_{1\leq t\leq T-1}\eta_{t})\sum_{t=1}^{T}\left|\eta_{t}^{-1}-\eta_{t-1}^{-1}\right|$
where (a) comes from the update rule (Eq. (2)), (b) is due to triangle inequality, and (c) follows from Lemma D.1. Thus, we have
$\sum_{t=1}^{T}\mathbb{E}_{S}\left[\nabla_{\hat{\tau}_{t}}l_{1-\alpha}(\hat{\tau}_{t},S)\right]\leq\frac{1}{1-\epsilon}\sqrt{2T\log\left(\frac{4}{\delta}\right)}+\frac{1+\underset{1\leq t\leq T-1}{\max}\eta_{t}}{1-\epsilon}\sum_{t=1}^{T}\left|\eta_{t}^{-1}-\eta_{t-1}^{-1}\right|+\frac{\epsilon}{1-\epsilon}\sum_{t=1}^{T}\left((1-\alpha)-\frac{1}{K}\mathbb{E}\left[\mathcal{C}_{t}(X_{t})\right]\right)$
By taking union bound and combining two parts, we can obtain that
$\quad\sum_{t=1}^{T}\nabla_{\tau_{t}}l_{1-\alpha}\left(\tau_{t},S_{t}\right)$ $\leq\sqrt{2T\log\left(\frac{4}{\delta}\right)}+\frac{1}{1-\epsilon}\sqrt{2T\log\left(\frac{4}{\delta}\right)}+\frac{1+\underset{1\leq t\leq T-1}{\max}\eta_{t}}{1-\epsilon}\sum_{t=1}^{T}\left|\eta_{t}^{-1}-\eta_{t-1}^{-1}\right|+$
$\quad\frac{\epsilon}{1-\epsilon}\sum_{t=1}^{T}\left(\frac{1}{K}\mathbb{E}\left[\mathcal{C}_{t}(X_{t})\right]-(1-\alpha)\right)$
$=\frac{2-\epsilon}{1-\epsilon}\sqrt{2T\log\left(\frac{4}{\delta}\right)}+\frac{1+\underset{1\leq t\leq T-1}{\max}\eta_{t}}{1-\epsilon}\sum_{t=1}^{T}\left|\eta_{t}^{-1}-\eta_{t-1}^{-1}\right|+\frac{\epsilon}{1-\epsilon}\sum_{t=1}^{T}\left((1-\alpha)-\frac{1}{K}\mathbb{E}\left[\mathcal{C}_{t}(X_{t})\right]\right)$
holds with at least $1-\delta$ probability. Recall that
We can conclude that
$\alpha-\frac{1}{T}\sum_{t=1}^{T}\mathds{1}\left\{Y_{t}\notin\mathcal{C}_{t}(X_{t})\right\}\leq$ $\frac{2-\epsilon}{1-\epsilon}\sqrt{2\log\left(\frac{4}{\delta}\right)}\cdot\frac{1}{\sqrt{T}}+\frac{1+\underset{1\leq t\leq T-1}{\max}\eta_{t}}{1-\epsilon}\sum_{t=1}^{T}\left|\eta_{t}^{-1}-\eta_{t-1}^{-1}\right|\cdot\frac{1}{T}+$
$\frac{\epsilon}{1-\epsilon}\cdot\frac{1}{T}\sum_{t=1}^{T}\left((1-\alpha)-\frac{1}{K}\mathbb{E}\left[\mathcal{C}_{t}(X_{t})\right]\right)$
D.2 Proof for Proposition 4.3
Proposition D.3 (Restatement of Proposition 4.3).
@@ Appendix D Proof for Section 4
=$\overset{(a)}{=}\left|\sum_{t=1}^{T}(\eta_{t}^{-1}-\eta_{t-1}^{-1})\cdot(\hat{\tau}_{T}-\hat{\tau}_{t})\right|$
@@
where (a) comes from the update rule (Eq. (6)), (b) is due to triangle inequality, and (c) follows from Lemma D.1. In addition, Lemma G.5 gives us that
@@
D.3 Proof for Proposition 4.4
Proposition D.4 (Restatement of Proposition 4.4).
@@
Lemma G.3 gives us that
@@
holds with at least $21-\delta/$ 2 probability. In addition, recall that
@@
Following from Lemma G.2, we have
@@
Lemma G.4 gives us that
@@
Appendix F Proof for Appendix B.2
Lemma F.1.
@@
We employ the proof technique from Lemma C.4 in Bhatnagar et al. (2023). The Assumption B.2 gives us that
@@
By Theorem 2.7 of Steinwart & Christmann (2011), we have
@@
Proposition F.2 (Restatement of Proposition B.3).
@@
By taking expectation condition on $(X_{t},Y_{t})$ (or equivalently on $\tilde{S}_{t}$ and $\{S_{t,y}\}_{y=1}^{K}$), and applying Lemma G.2 and Proposition 3.2, we have
@@
Continuing from Lemma F.1, we can conclude that
@@
Appendix G Helpful lemmas
Lemma G.1.
@@ Appendix G Helpful lemmas
$\mathbb{P}\left\{\tilde{S}=s\right\}$ =$=\mathbb{P}\left\{\tilde{S}=s|\tilde{S}=S\right\}\cdot\mathbb{P}\left\{\tilde{S}=S\right\}+\mathbb{P}\left\{\tilde{S}=s|\tilde{S}=\bar{S}\right\}\cdot\mathbb{P}\left\{\tilde{S}=\bar{S}\right\}$
=$=\mathbb{P}\left\{S=s\right\}\cdot\mathbb{P}\left\{\tilde{Y}=Y\right\}+\mathbb{P}\left\{\bar{S}=s\right\}\cdot\mathbb{P}\left\{\tilde{Y}=\bar{Y}\right\}$
@@
$\mathbb{P}\left\{\tilde{S}\leq s\right\}$ =$=\mathbb{P}\left\{\tilde{S}\leq s|\tilde{S}=S\right\}\cdot\mathbb{P}\left\{\tilde{S}=S\right\}+\mathbb{P}\left\{\tilde{S}\leq s|\tilde{S}=\bar{S}\right\}\cdot\mathbb{P}\left\{\tilde{S}=\bar{S}\right\}$
=$=\mathbb{P}\left\{S\leq s\right\}\cdot\mathbb{P}\left\{\tilde{Y}=Y\right\}+\mathbb{P}\left\{\bar{S}\leq s\right\}\cdot\mathbb{P}\left\{\tilde{Y}=\bar{Y}\right\}$
@@
$\mathbb{P}\left\{\tilde{S}>s\right\}$ =$=\mathbb{P}\left\{\tilde{S}>s|\tilde{S}=S\right\}\cdot\mathbb{P}\left\{\tilde{S}=S\right\}+\mathbb{P}\left\{\tilde{S}>s|\tilde{S}=\bar{S}\right\}\cdot\mathbb{P}\left\{\tilde{S}=\bar{S}\right\}$
=$=\mathbb{P}\left\{S>s\right\}\cdot\mathbb{P}\left\{\tilde{Y}=Y\right\}+\mathbb{P}\left\{\bar{S}>s\right\}\cdot\mathbb{P}\left\{\tilde{Y}=\bar{Y}\right\}$
@@
$-\underset{1\leq t\leq T-1}{\max}\eta_{t}\cdot\left(\alpha+\frac{\epsilon}{1-\epsilon}\right)\leq\hat{\tau}_{T}\leq 1+\underset{1\leq t\leq T-1}{\max}\eta_{t}\cdot\left(\frac{1}{1-\epsilon}-\alpha\right)$ (14)
@@
and Lemma G.2 gives us that
@@
$\mathbb{P}\left\{\left|\sum_{t=1}^{T}D_{t}\right|\leq t\right\}$ =$=\mathbb{P}\left\{\left|\sum_{t=1}^{T}\mathbb{E}_{\tilde{S}_{t},S_{t,y}}[\nabla_{\hat{\tau}_{t}}\tilde{l}_{1-\alpha}(\hat{\tau}_{t},\tilde{S}_{t},\{S_{t,y}\}_{y=1}^{K})]-\sum_{t=1}^{T}\nabla_{\hat{\tau}_{t}}\tilde{l}_{1-\alpha}(\hat{\tau}_{t},\tilde{S}_{t},\{S_{t,y}\}_{y=1}^{K})\right|\leq r\right\}$

### v3

@@ Appendix E Omitted proofs
E.5 Proof for Proposition 4.5
Proposition E.7 (Restatement of Proposition 4.5).
@@
$\overset{(a)}{=}\left|\sum_{t=1}^{T}(\eta_{t}^{-1}-\eta_{t-1}^{-1})\cdot(\hat{\tau}_{T}-\hat{\tau}_{t})\right|$
@@
where (a) comes from the update rule (Eq. (6)), (b) is due to triangle inequality, and (c) follows from Lemma LABEL:lemma:changing_lr_clean_hat_tau_bounded. In addition, Lemma F.5 gives us that
@@
E.6 Proof for Proposition 4.6
Proposition E.8 (Restatement of Proposition 4.6).
@@
Lemma F.3 gives us that
@@
holds with at least $1-\delta/2$ probability. In addition, recall that
@@
Following from Lemma F.2, we have
@@
Lemma F.4 gives us that
@@
E.8 Proof for Proposition B.3
Lemma E.10.
@@
We employ the proof technique from Lemma C.4 in [11]. The Assumption B.2 gives us that
@@
By Theorem 2.7 of [66], we have
@@
Proposition E.11 (Restatement of Proposition B.3).
@@
By taking expectation condition on $(X_{t},Y_{t})$ (or equivalently on $\tilde{S}_{t}$ and $\{S_{t,y}\}_{y=1}^{K}$), and applying Lemma F.2 and Proposition 4.1, we have
@@
Continuing from Lemma E.10, we can conclude that
@@
Appendix F Helpful lemmas
Lemma F.1.
@@ Appendix F Helpful lemmas
$\mathbb{P}\left\{\tilde{S}=s\right\}$ $=\mathbb{P}\left\{\tilde{S}=s|\tilde{S}=S\right\}\cdot\mathbb{P}\left\{\tilde{S}=S\right\}+\mathbb{P}\left\{\tilde{S}=s|\tilde{S}=\bar{S}\right\}\cdot\mathbb{P}\left\{\tilde{S}=\bar{S}\right\}$
$=\mathbb{P}\left\{S=s\right\}\cdot\mathbb{P}\left\{\tilde{Y}=Y\right\}+\mathbb{P}\left\{\bar{S}=s\right\}\cdot\mathbb{P}\left\{\tilde{Y}=\bar{Y}\right\}$
@@
$\mathbb{P}\left\{\tilde{S}\leq s\right\}$ $=\mathbb{P}\left\{\tilde{S}\leq s|\tilde{S}=S\right\}\cdot\mathbb{P}\left\{\tilde{S}=S\right\}+\mathbb{P}\left\{\tilde{S}\leq s|\tilde{S}=\bar{S}\right\}\cdot\mathbb{P}\left\{\tilde{S}=\bar{S}\right\}$
$=\mathbb{P}\left\{S\leq s\right\}\cdot\mathbb{P}\left\{\tilde{Y}=Y\right\}+\mathbb{P}\left\{\bar{S}\leq s\right\}\cdot\mathbb{P}\left\{\tilde{Y}=\bar{Y}\right\}$
@@
$\mathbb{P}\left\{\tilde{S}>s\right\}$ $=\mathbb{P}\left\{\tilde{S}>s|\tilde{S}=S\right\}\cdot\mathbb{P}\left\{\tilde{S}=S\right\}+\mathbb{P}\left\{\tilde{S}>s|\tilde{S}=\bar{S}\right\}\cdot\mathbb{P}\left\{\tilde{S}=\bar{S}\right\}$
$=\mathbb{P}\left\{S>s\right\}\cdot\mathbb{P}\left\{\tilde{Y}=Y\right\}+\mathbb{P}\left\{\bar{S}>s\right\}\cdot\mathbb{P}\left\{\tilde{Y}=\bar{Y}\right\}$
@@
$\underset{1\leq t\leq T-1}{\max}\eta_{t}\cdot\left(\alpha+\frac{\epsilon}{1-\epsilon}\right)\leq\hat{\tau}_{T}\leq 1+\underset{1\leq t\leq T-1}{\max}\eta_{t}\cdot\left(\frac{1}{1-\epsilon}-\alpha\right)$ (14)
@@
and Lemma F.2 gives us that
@@
$\mathbb{P}\left\{\left|\sum_{t=1}^{T}D_{t}\right|\leq t\right\}$ $=\mathbb{P}\left\{\left|\sum_{t=1}^{T}\mathbb{E}_{\tilde{S}_{t},S_{t,y}}[\nabla_{\hat{\tau}_{t}}\tilde{l}_{1-\alpha}(\hat{\tau}_{t},\tilde{S}_{t},\{S_{t,y}\}_{y=1}^{K})]-\sum_{t=1}^{T}\nabla_{\hat{\tau}_{t}}\tilde{l}_{1-\alpha}(\hat{\tau}_{t},\tilde{S}_{t},\{S_{t,y}\}_{y=1}^{K})\right|\leq r\right\}$